# Nonhermitian topological zero modes at smooth domain walls: Exact solutions

### v1

@@ Appendix B Exact solutions of the modified Jackiw-Rebbi equation
We found that the solutions $\varphi_{1,2}^{s}(x)$ can be written in closed form in terms of hypergeometric functions in the case where ${m}(y)$ and ${v}(y)$ can be expanded in powers of $y$, and these expansions can be truncated at the second order and first order, respectively, as
@@
where ${m}_{0}={m}_{L}$, ${m}_{1}={m}_{R}-{m}_{L}-{m}_{2}$, ${m}_{2}=2({m}_{L}+{m}_{R})-4{m}(0)$, ${v}_{0}={v}_{L}$, and ${v}_{1}={v}_{R}-{v}_{L}$. At the zeroth order, Eq. 46 describes the cases of constant fields ${m}(x)={m}_{L,R}={m}_{0}$ when ${m}_{1}={m}_{2}=0$ and ${v}(x)={v}_{L,R}={v}_{0}$ when ${v}_{1}=0$. At the first order, Eq. 46 describes the case where the fields follow an S-shaped curve interpolating between ${v}_{L}={v}_{0}$ and ${v}_{R}={v}_{0}+{v}_{1}$ for $x=\pm\infty$, and interpolating between ${m}_{L}={m}_{0}$ and ${m}_{R}={m}_{0}+{m}_{1}$ when ${m}_{2}=0$. In particular, a symmetric S-shaped curve interpolating between ${v}_{L}={v}_{0}$ and ${v}_{R}=-{v}_{0}$ for $x=\pm\infty$ is obtained if ${v}_{1}=-2{v}_{0}$. At the second order, ${m}(x)$ in Eq. 46 describes the symmetric Pöschl–Teller potential with ${m}_{L,R}={m}_{0}$ in the case where ${m}_{1}+{m}_{2}=0$ and an asymmetric Pöschl–Teller potential, i.e., a superposition of a Pöschl–Teller potential and an S-shaped potential otherwise.
If the expansion in Eq. 46 holds, Eq. 38 simplifies to
@@
±$\pm\frac{1}{2}\sqrt{\left(2ws{v}_{1}-1\right)^{2}+4w^{2}{m}_{2}},$ (48a)
@@
The explicit expressions for $F_{1,2}$ in Eqs. 42 and 40 are obtained by taking two solutions of the hypergeometric equation, i.e, choosing two linearly independent solutions out of the Kummer’s 24 solutions of the hypergeometric equation, which yields
@@
Notice that, if one considers higher order terms in Eq. 46, the resulting equation will not reduce to a hypergeometric equation in the general case. The equation above admits solutions that can be written in terms of two linearly independent solutions given by hypergeometric functions, although the choice is not unique. Hence, the explicit expressions for $F_{1,2}$ in Eq. 40 are obtained by taking
@@
where $a_{1},b_{1},c_{1}$ and $a_{2},b_{2},c_{2}$ are obtained from Eq. 48 by substituting $({\alpha_{L}},{\alpha_{R}})=(\alpha_{L}^{+},\alpha_{R}^{-})$ and $({\alpha_{L}},{\alpha_{R}})=(\alpha_{L}^{-},\alpha_{R}^{+})$, respectively, giving
$a_{1,2}$ =±$=\pm w({{q}_{L}}-{{q}_{R}})$
@@
provided that the corresponding solutions of the modified Jackiw-Rebbi equation in Eq. 40 are linearly independent and can satisfy the boundary conditions. We notice that $a_{1}-a_{2}=b_{1}-b_{2}=c_{1}-c_{2}$ and that $b_{1,2}=1-a_{2,1}$.
Since ${q}_{L}$ and ${q}_{R}$ cannot be negative real numbers by definition, the hypergeometric functions above are well-defined, i.e., the hypergeometric function converges since $\mathrm{Re}(c_{1,2})\geq 1$ and therefore $c_{1,2}\notin\mathbb{Z}_{\leq}$. Moreover, to satisfy the boundary conditions, we must ensure that the functions defined above are sufficiently well-behaved at the extremes of the interval $0\leq y\leq 1$. We have that ${{\,}_{2}F_{1}}(a,b,c,0)=1$ which gives $F_{1}(0)=F_{2}(1)=1$. Furthermore, the asymptotic behavior of ${{\,}_{2}F_{1}}(a,b,c,y\to 1)$ depends on the value of $c-a-b$. Since, $c_{1}-a_{1}-b_{1}=c_{2}-1=2{w{q}_{R}}=2w\sqrt{{{v}}_{R}^{2}+{m}_{R}}$, by using the asymptotic properties of the hypergeometric functions, one obtains the following asymptotic behaviors, assuming $c_{1}-a_{1},c_{1}-b_{1}\notin\mathbb{Z}_{\leq}$. If either $a_{1}\in\mathbb{Z}_{\leq}$ or $b_{1}\in\mathbb{Z}_{\leq}$, the hypergeometric function reduces to a polynomial and one has
@@
Analogously, if $c_{2}-a_{2},c_{2}-b_{2}\notin\mathbb{Z}_{\leq}$ and either $a_{1}\in\mathbb{Z}_{\leq}$ or $b_{1}\in\mathbb{Z}_{\leq}$, then $F_{2}(0)$ is finite and nonzero, while if $a_{1},b_{1},c_{2}-a_{2},c_{2}-b_{2}\notin\mathbb{Z}_{\leq}$ then $F_{2}(y)$ converges for $y\to 0$ to a nonzero value if $\mathrm{Re}({q}_{L})>0$ (e.g., ${{v}}_{L}^{2}+{m}_{L}>0$ in the asymptotically hermitian case ${v}_{L},{m}_{L}\in\mathbb{R}$), it is bounded with oscillating behavior if $\mathrm{Re}({q}_{L})=0$ and ${q}_{L}\neq 0$ (e.g., ${{v}}_{L}^{2}+{m}_{L}<0$ in the asymptotically hermitian case ${v}_{L},{m}_{L}\in\mathbb{R}$), and it diverges logarithmically if ${q}_{L}=0$ (i.e., ${{v}}_{L}^{2}+{m}_{L}=0$) with asymptotic behavior given by analogous equations as in Eqs. 52, 53, 54 and 55 where $a_{1},b_{1},c_{1}\to a_{2},b_{2},c_{2}$ and ${q}_{R}\to{q}_{L}$.
Hence, if $c_{1,2}-a_{1,2},c_{1,2}-b_{1,2}\notin\mathbb{Z}_{\leq}$, then $0<|F_{1,2}(y)|<\infty$ on the whole interval $0\leq y\leq 1$ or diverges logarithmically at the extremes of the interval. This ensures that $|\varphi_{1,2}^{s}(y\to 0)|\sim y^{w\mathrm{Re}(\alpha_{L}^{\pm}})$ and $|\varphi_{1,2}^{s}(y\to 1)|\sim(1-y)^{w\mathrm{Re}(\alpha_{R}^{\mp}})$. Finally, we notice that the choice of the functions $F_{1,2}$ in Eq. 50 ensures that the solutions of the modified Jackiw-Rebbi equation in Eq. 40 are linearly independent, since the Wronskian of the two solutions is
@@
Transforming back to the real line $-\infty<x<\infty$, the functions $F_{1,2}$ in Eq. 42 are given by
@@
One has that $F_{1,2}(x\to\mp\infty)=1$. Furthermore, if $c_{1,2}-a_{1,2},c_{1,2}-b_{1,2}\notin\mathbb{Z}_{\leq}$, then $F_{1}(x)$ converges for $x\to\infty$ to a nonzero value if $\mathrm{Re}({q}_{R})>0$ (e.g., ${{v}}_{R}^{2}+{m}_{R}>0$ in the asymptotically hermitian case ${v}_{R},{m}_{R}\in\mathbb{R}$) while $F_{2}(x)$ converges for $x\to-\infty$ to a nonzero value if $\mathrm{Re}({q}_{L})>0$ (e.g., ${{v}}_{L}^{2}+{m}_{L}>0$ in the asymptotically hermitian case ${v}_{L},{m}_{L}\in\mathbb{R}$); $F_{1}(x)$ is bounded with oscillating behavior if $\mathrm{Re}({q}_{R})=0$ and ${q}_{R}\neq 0$ (e.g., ${{v}}_{R}^{2}+{m}_{R}<0$ in the asymptotically hermitian case ${v}_{R},{m}_{R}\in\mathbb{R}$) while $F_{2}(x)$ is bounded with oscillating behavior if $\mathrm{Re}({q}_{L})=0$ and ${q}_{L}\neq 0$ (e.g., ${{v}}_{L}^{2}+{m}_{L}<0$ in the asymptotically hermitian case ${v}_{L},{m}_{L}\in\mathbb{R}$); $F_{1}(x)$ diverges polynomially if ${q}_{R}=0$ (i.e., ${{v}}_{R}^{2}+{m}_{R}=0$) while $F_{2}(x)$ diverges polynomially if ${q}_{L}=0$ (i.e., ${{v}}_{L}^{2}+{m}_{L}=0$). These asymptotic behaviors are summarized in Table 1. It follows that, if $c_{1,2}-a_{1,2},c_{1,2}-b_{1,2}\notin\mathbb{Z}_{\leq}$, then $0<|F_{1,2}(x)|<\infty$ on the whole real line or diverges polynomially for $x\to\pm\infty$.
As previously observed, in the hermitian case $\mathrm{Im}({m}(x))=\mathrm{Im}({v}(x))=0$, then the modified Jackiw-Rebbi equation in Eq. 3 is real, and hence the solutions $\varphi_{1,2}^{s}(x)$ are either real or can be linearly combined into a set of linearly independent solutions which are real. In the nonhermitian case when $\mathrm{Im}({m}(x)\neq 0$ or $\mathrm{Im}({v}(x))\neq 0$ instead, the solutions $\varphi_{1,2}^{s}(x)$ cannot be linearly combined into a set of linearly independent solutions which are real.
In Table I of Ref. marra_topological_2024 are shown some special cases where the analytical expression of the general solution simplifies. In particular, we consider the cases where ${v}(x)={v}_{L,R}={v}$ is constant or a symmetric S-shaped curve (${v}_{0}={v}$ and ${v}_{1}=-2{v}$) with ${v}_{L,R}=\pm{v}$, and the cases where ${m}(x)={m}$ is constant or a symmetric Pöschl–Teller potential (${m}_{0}={m}$ and ${m}_{1}+{m}_{2}=0$), with ${m}_{L,R}={m}$, and more generally the case where ${q}_{L}={q}_{R}={q}$. In the cases considered, we have $|{v}_{L,R}|={v}$, and thus, we can assume ${v}\in\mathbb{R}$ up to a gauge transformation.
@@
The correspondence between the hypergeometric functions and associated Legendre functions is the consequence of the fact that the hypergeometric equation can be transformed into a Legendre’s differential equation by mapping the interval $0\leq y\leq 1$ to the interval $-1\leq z\leq 1$ by the substitution $z=1-2y=\tanh(x/2w)$.

### v2

@@ Appendix B Exact solutions of the modified Jackiw-Rebbi equation
We found that the solutions $\varphi_{1,2}^{s}(x)$ can be written in closed form in terms of hypergeometric functions in the case where m(y){m}(y) and v(y){v}(y) can be expanded in powers of $y$, and these expansions can be truncated at the second order and first order, respectively, as
@@
where ${m}_{0}={m}_{L}$, ${m}_{1}={m}_{R}-{m}_{L}-{m}_{2}$, ${m}_{2}=2({m}_{L}+{m}_{R})-4{m}(0)$, ${v}_{0}={v}_{L}$, and ${v}_{1}={v}_{R}-{v}_{L}$. At the zeroth order, Eq.˜46 describes the cases of constant fields ${m}(x)={m}_{L,R}={m}_{0}$ when ${m}_{1}={m}_{2}=0$ and ${v}(x)={v}_{L,R}={v}_{0}$ when ${v}_{1}=0$. At the first order, Eq.˜46 describes the case where the fields follow an S-shaped curve interpolating between ${v}_{L}={v}_{0}$ and ${v}_{R}={v}_{0}+{v}_{1}$ for $x=\pm\infty$, and interpolating between ${m}_{L}={m}_{0}$ and ${m}_{R}={m}_{0}+{m}_{1}$ when ${m}_{2}=0$. In particular, a symmetric S-shaped curve interpolating between ${v}_{L}={v}_{0}$ and ${v}_{R}=-{v}_{0}$ for $x=\pm\infty$ is obtained if ${v}_{1}=-2{v}_{0}$. At the second order, m(x){m}(x) in Eq.˜46 describes the symmetric Pöschl–Teller potential with ${m}_{L,R}={m}_{0}$ in the case where ${m}_{1}+{m}_{2}=0$ and an asymmetric Pöschl–Teller potential, i.e., a superposition of a Pöschl–Teller potential and an S-shaped potential otherwise.
If the expansion in Eq.˜46 holds, Eq.˜38 simplifies to
@@
$\pm\frac{1}{2}\sqrt{\left(2ws{v}_{1}-1\right)^{2}+4w^{2}{m}_{2}},$ (48a)
@@
The explicit expressions for $F_{1,2}$ in Eqs.˜42 and 40 are obtained by taking two solutions of the hypergeometric equation, i.e, choosing two linearly independent solutions out of the Kummer’s 24 solutions of the hypergeometric equation, which yields
@@
Notice that, if one considers higher order terms in Eq.˜46, the resulting equation will not reduce to a hypergeometric equation in the general case. The equation above admits solutions that can be written in terms of two linearly independent solutions given by hypergeometric functions, although the choice is not unique. Hence, the explicit expressions for $F_{1,2}$ in Eq.˜40 are obtained by taking
@@
where $a_{1},b_{1},c_{1}$ and $a_{2},b_{2},c_{2}$ are obtained from Eq.˜48 by substituting $({\alpha_{L}},{\alpha_{R}})=(\alpha_{L}^{+},\alpha_{R}^{-})$ and $({\alpha_{L}},{\alpha_{R}})=(\alpha_{L}^{-},\alpha_{R}^{+})$, respectively, giving
$a_{1,2}$ $=\pm w({{q}_{L}}-{{q}_{R}})$
@@
provided that the corresponding solutions of the modified Jackiw-Rebbi equation in Eq.˜40 are linearly independent and can satisfy the boundary conditions. We notice that $a_{1}-a_{2}=b_{1}-b_{2}=c_{1}-c_{2}$ and that $b_{1,2}=1-a_{2,1}$.
Since ${q}_{L}$ and ${q}_{R}$ cannot be negative real numbers by definition, the hypergeometric functions above are well-defined, i.e., the hypergeometric function converges since $\mathrm{Re}(c_{1,2})\geq 1$ and therefore $c_{1,2}\notin\mathbb{Z}_{\leq}$. Moreover, to satisfy the boundary conditions, we must ensure that the functions defined above are sufficiently well-behaved at the extremes of the interval $0\leq y\leq 1$. We have that ${{\,}_{2}F_{1}}(a,b,c,0)=1$ which gives $F_{1}(0)=F_{2}(1)=1$. Furthermore, the asymptotic behavior of ${{\,}_{2}F_{1}}(a,b,c,y\to 1)$ depends on the value of cabc-a-b. Since, $c_{1}-a_{1}-b_{1}=c_{2}-1=2{w{q}_{R}}=2w\sqrt{{{v}}_{R}^{2}+{m}_{R}}$, by using the asymptotic properties of the hypergeometric functions, one obtains the following asymptotic behaviors, assuming $c_{1}-a_{1},c_{1}-b_{1}\notin\mathbb{Z}_{\leq}$. If either $a_{1}\in\mathbb{Z}_{\leq}$ or $b_{1}\in\mathbb{Z}_{\leq}$, the hypergeometric function reduces to a polynomial and one has
@@
Analogously, if $c_{2}-a_{2},c_{2}-b_{2}\notin\mathbb{Z}_{\leq}$ and either $a_{1}\in\mathbb{Z}_{\leq}$ or $b_{1}\in\mathbb{Z}_{\leq}$, then $F_{2}(0)$ is finite and nonzero, while if $a_{1},b_{1},c_{2}-a_{2},c_{2}-b_{2}\notin\mathbb{Z}_{\leq}$ then $F_{2}(y)$ converges for $y\to 0$ to a nonzero value if $\mathrm{Re}({q}_{L})>0$ (e.g., ${{v}}_{L}^{2}+{m}_{L}>0$ in the asymptotically hermitian case ${v}_{L},{m}_{L}\in\mathbb{R}$), it is bounded with oscillating behavior if $\mathrm{Re}({q}_{L})=0$ and ${q}_{L}\neq 0$ (e.g., ${{v}}_{L}^{2}+{m}_{L}<0$ in the asymptotically hermitian case ${v}_{L},{m}_{L}\in\mathbb{R}$), and it diverges logarithmically if ${q}_{L}=0$ (i.e., ${{v}}_{L}^{2}+{m}_{L}=0$) with asymptotic behavior given by analogous equations as in Eqs.˜52, 53, 54 and 55 where $a_{1},b_{1},c_{1}\to a_{2},b_{2},c_{2}$ and ${q}_{R}\to{q}_{L}$.
Hence, if $c_{1,2}-a_{1,2},c_{1,2}-b_{1,2}\notin\mathbb{Z}_{\leq}$, then $0<|F_{1,2}(y)|<\infty$ on the whole interval $0\leq y\leq 1$ or diverges logarithmically at the extremes of the interval. This ensures that $|\varphi_{1,2}^{s}(y\to 0)|\sim y^{w\mathrm{Re}(\alpha_{L}^{\pm})}$ and $|\varphi_{1,2}^{s}(y\to 1)|\sim(1-y)^{w\mathrm{Re}(\alpha_{R}^{\mp})}$. Finally, we notice that the choice of the functions $F_{1,2}$ in Eq.˜50 ensures that the solutions of the modified Jackiw-Rebbi equation in Eq.˜40 are linearly independent, since the Wronskian of the two solutions is
@@
Transforming back to the real line $(-\infty,\infty)$, the functions $F_{1,2}$ in Eq.˜42 are given by
@@
One has that $F_{1,2}(x\to\mp\infty)=1$. Furthermore, if $c_{1,2}-a_{1,2},c_{1,2}-b_{1,2}\notin\mathbb{Z}_{\leq}$, then $F_{1}(x)$ converges for $x\to\infty$ to a nonzero value if $\mathrm{Re}({q}_{R})>0$ (e.g., ${{v}}_{R}^{2}+{m}_{R}>0$ in the asymptotically hermitian case ${v}_{R},{m}_{R}\in\mathbb{R}$) while $F_{2}(x)$ converges for $x\to$-$\infty$ to a nonzero value if $\mathrm{Re}({q}_{L})>0$ (e.g., ${{v}}_{L}^{2}+{m}_{L}>0$ in the asymptotically hermitian case ${v}_{L},{m}_{L}\in\mathbb{R}$); $F_{1}(x)$ is bounded with oscillating behavior if $\mathrm{Re}({q}_{R})=0$ and ${q}_{R}\neq 0$ (e.g., ${{v}}_{R}^{2}+{m}_{R}<0$ in the asymptotically hermitian case ${v}_{R},{m}_{R}\in\mathbb{R}$) while $F_{2}(x)$ is bounded with oscillating behavior if $\mathrm{Re}({q}_{L})=0$ and ${q}_{L}\neq 0$ (e.g., ${{v}}_{L}^{2}+{m}_{L}<0$ in the asymptotically hermitian case ${v}_{L},{m}_{L}\in\mathbb{R}$); $F_{1}(x)$ diverges polynomially if ${q}_{R}=0$ (i.e., ${{v}}_{R}^{2}+{m}_{R}=0$) while $F_{2}(x)$ diverges polynomially if ${q}_{L}=0$ (i.e., ${{v}}_{L}^{2}+{m}_{L}=0$). These asymptotic behaviors are summarized in Table˜1. It follows that, if $c_{1,2}-a_{1,2},c_{1,2}-b_{1,2}\notin\mathbb{Z}_{\leq}$, then $0<|F_{1,2}(x)|<\infty$ on the whole real line or diverges polynomially for $x\to\pm\infty$.
As previously observed, in the hermitian case $\mathrm{Im}({m}(x))=\mathrm{Im}({v}(x))=0$, then the modified Jackiw-Rebbi equation in Eq.˜2 is real, and hence the solutions $\varphi_{1,2}^{s}(x)$ are either real or can be linearly combined into a set of linearly independent solutions which are real. In the nonhermitian case when $\mathrm{Im}({m}(x)\neq 0$ or $\mathrm{Im}({v}(x))\neq 0$ instead, the solutions $\varphi_{1,2}^{s}(x)$ cannot be linearly combined into a set of linearly independent solutions which are real.
In Table I of Ref. marra_topological_2024 are shown some special cases where the analytical expression of the general solution simplifies. In particular, we consider the cases where ${v}(x)={v}_{L,R}={v}$ is constant or a symmetric S-shaped curve (${v}_{0}={v}$ and ${v}_{1}=-2{v}$) with ${v}_{L,R}=\pm{v}$, and the cases where m(x)=m{m}(x)={m} is constant or a symmetric Pöschl–Teller potential (${m}_{0}={m}$ and ${m}_{1}+{m}_{2}=0$), with ${m}_{L,R}={m}$, and more generally the case where ${q}_{L}={q}_{R}={q}$. In the cases considered, we have $|{v}_{L,R}|={v}$, and thus, we can assume ${v}\in\mathbb{R}$ up to a gauge transformation.
@@
The correspondence between the hypergeometric functions and associated Legendre functions is the consequence of the fact that the hypergeometric equation can be transformed into a Legendre’s differential equation by mapping the interval $0\leq y\leq 1$ to the interval $1\leq z\leq 1$ by the substitution $z=1-2y=\tanh(x/2w)$.